# Turbocharging Gaussian Process Inference with Approximate Sketch-and-Project

### v1

@@ Appendix C Additional Algorithmic Details for ADASAP
We provide additional implementation details for ADASAP. Section C.1 describes how we distribute matrix-matrix products across rows and columns in the algorithms ColDistMatMat and RowDistMatMat. Section C.2 describes the practical implementation of the randomized Nyström approximation and provides pseudocode for RandNysAppx. Section C.3 describes how we compute preconditioned smoothness constants and provides pseuodocode for GetStepsize. Section C.4 provides pseudocode for Nesterov acceleration in NestAcc. Section C.5 provides a scaling plot illustrating the speedups achieved by using multiple GPUs in ADASAP.
@@ C.1 Distributed matrix-matrix products
Here, we provide details for how we implement the distributed matrix-matrix products in ADASAP. ColDistMatMat (Algorithm 3) shows how we distribute the matrix-matrix product $K_{\mathcal{B}n}W_{t}$ in ADASAP and RowDistMatMat (Algorithm 4) shows how we distribute the calculation of the Nyström sketch in ADASAP. Our implementation of these algorithms uses torch.multiprocessing to spawn a CUDA context on each device (i.e., a worker) and uses pykeops to generate the column and row block oracles.
@@
$(K_{\mathcal{B}n}W)_{i}\leftarrow\mathcal{K}^{\textrm{col}}_{\mathcal{W}_{i}}[W_{i}]$triangleright Compute column block products in parallel
@@
$(K_{\mathcal{B}\mathcal{B}}\Omega)_{i}\leftarrow\mathcal{K}^{\textrm{row}}_{\mathcal{W}_{i}}[\Omega]$triangleright Compute row block products in parallel
@@ C.2 Randomized Nyström approximation
Here, we present a practical implementation of the Nyström approximation used in ADASAP (Algorithm 2) in RandNysAppx (Algorithm 5). The Randomize Nyström approximation of $K_{\mathcal{B}\mathcal{B}}$ with test matrix $\Omega\in\mathbb{R}^{b\times r}$ [Tropp et al., 2017] is given by:
@@
The preceding formula is numerically unstable, so ADASAP uses RandNysAppx, which is based on Tropp et al. [2017, Algorithm 3]. $\text{eps}(x)$ is defined as the positive distance between $x$ and the next largest floating point number of the same precision as $x$. The resulting Nyström approximation $\hat{M}$ is given by $USU^{T}$, where $U\in\mathbb{R}^{p\times r}$ is an orthogonal matrix that contains the approximate top-$r$ eigenvectors of $M$, and $S\in\mathbb{R}^{r}$ contains the top-$r$ eigenvalues of $M$. The Nyström approximation is positive-semidefinite but may have eigenvalues that are equal to 00. In our algorithms, this approximation is always used in conjunction with a regularizer to ensure positive definiteness.
@@
$\Delta\leftarrow\mathrm{eps}(\Omega^{T}M\Omega.\mathrm{dtype})\cdot\operatorname{Tr}(\Omega^{T}M\Omega)$triangleright Compute shift for stability
$C\leftarrow\text{chol}(\Omega^{T}M\Omega+\Delta\Omega^{T}\Omega)$triangleright Cholesky decomposition: $C^{T}C=\Omega^{T}M\Omega+\Delta\Omega^{T}\Omega$
$B\leftarrow YC^{-1}$triangleright Triangular solve
$[U,\Sigma,\sim]\leftarrow\text{svd}(B,0)$triangleright Thin SVD
$S\leftarrow\text{max}\{0,\operatorname{diag}(S^{2}-\Delta I)\}$triangleright Compute eigs, and remove shift with element-wise max
@@
The dominant cost in Algorithm 5 is computing the SVD of $B$ at a cost of $\mathcal{O}(pr^{2})$. This is the source of the $\mathcal{O}(pr^{2})$ cost in Phase III of ADASAP.
@@ C.2.1 Applying the Nyström approximation to a vector
We also use the randomized Nyström approximation to compute preconditioned smoothness constants in GetStepsize (Algorithm 6). This computation requires the calculation $(P+\rho I)^{-1/2}v$ for some $v\in\mathbb{R}^{p}$, which can also be performed in $\mathcal{O}(pr)$ time using the Woodbury formula:
@@
Unlike Eq. 8, we find using (9) in GetStepsize yields excellent performance, i.e., we do not need to perform any additional stabilization.
@@ C.3 Computing the stepsize
Here, we provide the details of the GetStepsize procedure in ADASAP (Algorithm 2) for automatically computing the stepsize. Our procedure is inspired by Rathore et al. [2025], who show that approximate SAP with the Nyström approximation converges when
@@
using only matrix-vector products with the matrices $H$ and $(P+\rho I)^{-1/2}$. When $P$ is calculated using RandNysAppx, GetStepsize can efficiently compute a matrix-vector product with $(P+\rho I)^{-1/2}$ using (9). In practice, we find that 10 iterations of randomized powering are sufficient for estimating the preconditioned smoothness constant. The pseudocode for randomized powering, based on the presentation in Martinsson and Tropp [2020], is shown in Algorithm 6. Since ADASAP only runs GetStepsize for $10$ iterations, the total cost of the procedure is $\mathcal{O}(b^{2})$, which makes it the source of the $\mathcal{O}(b^{2})$ cost of Phase III in Algorithm 2.
@@
$v_{0}\leftarrow v_{0}/\|v_{0}\|_{2}$triangleright Normalize
@@
     $v_{i+1}\leftarrow v_{i+1}/\|v_{i+1}\|_{2}$triangleright Normalize
@@ C.4 Nesterov acceleration
We present pseudocode for Nesterov acceleration in NestAcc (Algorithm 7).
@@ C.5 Parallel scaling of ADASAP
Here we present Fig. 4, which shows the parallel scaling of ADASAP on the taxi dataset.
@@ D.2 Optimizer hyperparameters
For GP inference on large-scale datasets, we use blocksize $b=n/100$ in ADASAP, ADASAP-I, and SDD; blocksize $b=n/2{,}000$ for transporation data analysis, and $b=n/5$ for Bayesian optimization,
@@ D.5 Additional details for Bayesian optimization
We draw the target functions $f:[0,1]^{8}\rightarrow\mathbb{R}$ from a zero-mean GP prior with Matérn-3/2 kernel using 5,000 random features. At each iteration, we choose the acquisition points using parallel Thompson sampling [Hernández-Lobato et al., 2017]. As part of this process, we use the multi-start gradient optimization maximization strategy given in Lin et al. [2023]. At each step, we acquire 1,000 new points, which we use to evaluate the objective function. Concretely, if $x_{\mathrm{new}}$ is an acquired point, we compute $y_{\mathrm{new}}=f(x_{\mathrm{new}})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,10^{-6})$. We then add $(x_{\mathrm{new}},y_{\mathrm{new}})$ to the training data for the next step of optimization. In our experiments, we initialize all methods with a dataset consisting of 250,000 observations sampled uniformly at random from $[0,1]^{8}$.
D.6 Additional timing plots for Section 5.1
Here we present timing plots for the datasets used in Section 5.1 that were not shown in the main paper. Fig. 5 shows that all the methods (except PCG) perform similarly on both test RMSE and test mean NLL. However, Fig. 6 shows that ADASAP and PCG attain a much lower train RMSE than the competitors. This suggests that the similar performance of the methods on test RMSE and test mean NLL is not due to differences in optimization, but rather, it is because the datasets are not well-modeled by Gaussian processes.

### v2

@@ Appendix C Additional Algorithmic Details for ADASAP
We provide additional implementation details for ADASAP. Section˜C.1 describes how we distribute matrix-matrix products across rows and columns in the algorithms ColDistMatMat and RowDistMatMat. Section˜C.2 describes the practical implementation of the randomized Nyström approximation and provides pseudocode for RandNysAppx. Section˜C.3 describes how we compute preconditioned smoothness constants and provides pseuodocode for GetStepsize. Section˜C.4 provides pseudocode for Nesterov acceleration in NestAcc. Section˜C.5 provides a scaling plot illustrating the speedups achieved by using multiple GPUs in ADASAP. Section˜C.6 investigates the impact of tail averaging on the performance of ADASAP.
@@ C.1 Distributed matrix-matrix products
Here, we provide details for how we implement the distributed matrix-matrix products in ADASAP. ColDistMatMat (Algorithm˜3) shows how we distribute the matrix-matrix product $K_{\mathcal{B}n}W_{t}$ in ADASAP and RowDistMatMat (Algorithm˜4) shows how we distribute the calculation of the Nyström sketch in ADASAP. Our implementation of these algorithms uses torch.multiprocessing to spawn a CUDA context on each device (i.e., a worker) and uses pykeops to generate the column and row block oracles.
@@
$(K_{\mathcal{B}n}W)_{i}\leftarrow\mathcal{K}^{\textrm{col}}_{\mathcal{W}_{i}}[W_{i}]$ $\triangleright$ Compute column block products in parallel
@@
$(K_{\mathcal{B}\mathcal{B}}\Omega)_{i}\leftarrow\mathcal{K}^{\textrm{row}}_{\mathcal{W}_{i}}[\Omega]$ $\triangleright$ Compute row block products in parallel
@@ C.2 Randomized Nyström approximation
Here, we present a practical implementation of the Nyström approximation used in ADASAP (Algorithm˜2) in RandNysAppx (Algorithm˜5). The Randomized Nyström approximation of $K_{\mathcal{B}\mathcal{B}}$ with test matrix $\Omega\in\mathbb{R}^{b\times r}$ [Tropp et al., 2017] is given by:
@@
The preceding formula is numerically unstable, so ADASAP uses RandNysAppx, which is based on Tropp et al. [2017, Algorithm 3]. $\text{eps}(x)$ is defined as the positive distance between $x$ and the next largest floating point number of the same precision as $x$. The resulting Nyström approximation $\hat{M}$ is given by $USU^{T}$, where $U\in\mathbb{R}^{p\times r}$ is an orthogonal matrix that contains the approximate top-$r$ eigenvectors of $M$, and $S\in\mathbb{R}^{r}$ contains the top-$r$ eigenvalues of $M$. The Nyström approximation is positive-semidefinite but may have eigenvalues that are equal to 0. In our algorithms, this approximation is always used in conjunction with a regularizer to ensure positive definiteness.
@@
$\Delta\leftarrow\mathrm{eps}(\Omega^{T}M\Omega.\mathrm{dtype})\cdot\operatorname{Tr}(\Omega^{T}M\Omega)$ $\triangleright$ Compute shift for stability
$C\leftarrow\text{chol}(\Omega^{T}M\Omega+\Delta\Omega^{T}\Omega)$ $\triangleright$ Cholesky decomposition: $C^{T}C=\Omega^{T}M\Omega+\Delta\Omega^{T}\Omega$
$B\leftarrow YC^{-1}$ $\triangleright$ Triangular solve
$U,\Sigma,\sim]\leftarrow\text{svd}(B,0) \triangleright$ Thin SVD
$S\leftarrow\text{max}\{0,\operatorname{diag}(S^{2}-\Delta I)\}$ $\triangleright$ Compute eigs, and remove shift with element-wise max
@@
The dominant cost in Algorithm˜5 is computing the SVD of $B$ at a cost of $\mathcal{O}(pr^{2})$. This is the source of the $\mathcal{O}(pr^{2})$ cost in Phase III of ADASAP.
@@ C.2.1 Applying the Nyström approximation to a vector
We also use the randomized Nyström approximation to compute preconditioned smoothness constants in GetStepsize (Algorithm˜6). This computation requires the calculation $(P+\rho I)^{-1/2}v$ for some $v\in\mathbb{R}^{p}$, which can also be performed in $\mathcal{O}(pr)$ time using the Woodbury formula:
@@
Unlike Eq.˜8, we find using (9) in GetStepsize yields excellent performance, i.e., we do not need to perform any additional stabilization.
@@ C.3 Computing the stepsize
Here, we provide the details of the GetStepsize procedure in ADASAP (Algorithm˜2) for automatically computing the stepsize. Our procedure is inspired by Rathore et al. [2025], who show that approximate SAP with the Nyström approximation converges when
@@
using only matrix-vector products with the matrices $H$ and $(P+\rho I)^{-1/2}$. When $P$ is calculated using RandNysAppx, GetStepsize can efficiently compute a matrix-vector product with $(P+\rho I)^{-1/2}$ using (9). In practice, we find that 10 iterations of randomized powering are sufficient for estimating the preconditioned smoothness constant. The pseudocode for randomized powering, based on the presentation in Martinsson and Tropp [2020], is shown in Algorithm˜6. Since ADASAP only runs GetStepsize for $10$ iterations, the total cost of the procedure is $\mathcal{O}(b^{2})$, which makes it the source of the $\mathcal{O}(b^{2})$ cost of Phase III in Algorithm˜2.
@@
$v_{0}\leftarrow v_{0}/\|v_{0}\|_{2}$ $\triangleright$ Normalize
@@
  $v_{i+1}\leftarrow v_{i+1}/\|v_{i+1}\|_{2}$ $\triangleright$ Normalize
@@ C.4 Nesterov acceleration
We present pseudocode for Nesterov acceleration in NestAcc (Algorithm˜7).
@@ C.5 Parallel scaling of ADASAP
Here we present Fig.˜4, which shows the parallel scaling of ADASAP on the taxi dataset.
@@
C.6 Impact of tail averaging on performance
The theoretical results in Section˜3 require tail averaging. Here we assess whether tail averaging leads to practical improvements in the convergence of sketch-and-project algorithms for kernel ridge regression. To do so, we run a synthetic experiment using an RBF kernel with $n=1000$ samples.
Fig.˜5 displays the relative errors over the top-$\ell$ subspace, which are computed using the expression
$\|\mathrm{proj}_{\ell}(\hat{m})-\mathrm{proj}_{\ell}(m_{n})\|^{2}_{\mathcal{H}}/\|\mathrm{proj}_{\ell}(m_{n})\|^{2}_{\mathcal{H}}.$
When $\ell=10$ and the blocksize $b$ is small, tail averaging slightly improves the convergence rate. However, when $\ell\in\{100,1000\}$, tail averaging does not improve the convergence rate. In fact, as the blocksize increases, tail averaging results in slower convergence!
Figure 5: Performance of ADASAP with and without tail averaging. One “data pass” corresponds to one pass through the kernel matrix. Tail averaging does not improve convergence by a substantial margin.
@@ D.2 Optimizer hyperparameters
For GP inference on large-scale datasets, we use blocksize $b=n/100$ in ADASAP, ADASAP-I, and SDD; blocksize b=n/2,000b=n/2{,}000 for transporation data analysis, and $b=n/5$ for Bayesian optimization,
@@ D.5 Additional details for Bayesian optimization
We draw the target functions $f:[0,1]^{8}\rightarrow\mathbb{R}$ from a zero-mean GP prior with Matérn-3/2 kernel using 5,000 random features. At each iteration, we choose the acquisition points using parallel Thompson sampling [Hernández-Lobato et al., 2017]. As part of this process, we use the multi-start gradient optimization maximization strategy given in Lin et al. [2023]. At each step, we acquire 1,000 new points, which we use to evaluate the objective function. Concretely, if $x_{\mathrm{new}}$ is an acquired point, we compute $y_{\mathrm{new}}=f(x_{\mathrm{new}})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,10^{-6})$. We then add $(x_{\mathrm{new}},y_{\mathrm{new}})$ to the training data for the next step of optimization. In our experiments, we initialize all methods with a dataset consisting of 250,000 observations sampled uniformly at random from $0,1]^{8}$.
D.6 Additional timing plots for Section˜5.1
Here we present timing plots for the datasets used in Section˜5.1 that were not shown in the main paper. Fig.˜6 shows that all the methods (except PCG) perform similarly on both test RMSE and test mean NLL. However, Fig.˜7 shows that ADASAP and PCG attain a much lower train RMSE than the competitors. This suggests that the similar performance of the methods on test RMSE and test mean NLL is not due to differences in optimization, but rather, it is because the datasets are not well-modeled by Gaussian processes.